# CATP-LLM: Empowering Large Language Models for Cost-Aware Tool Planning

### v2

@@ Abstract
Utilizing large language models (LLMs) for tool planning has emerged as a promising avenue for developing general AI systems, where LLMs automatically schedule external tools (e.g. vision models) to tackle complex tasks based on task descriptions. To push this paradigm toward practical applications, it is crucial for LLMs to consider tool execution costs (e.g. execution time) for tool planning. Unfortunately, prior studies overlook the tool execution costs, leading to the generation of expensive plans of which the costs outweigh task performance. To fill this gap, we propose the Cost-Aware Tool Planning with LLMs (CATP-LLM) framework, which for the first time provides a coherent design to empower LLMs for cost-aware tool planning. Specifically, CATP-LLM incorporates a tool planning language to enhance the LLM to generate non-sequential plans of multiple branches for efficient concurrent tool execution and cost reduction. Moreover, it further designs a cost-aware offline reinforcement learning algorithm to fine-tune the LLM to optimize the performance-cost trade-off in tool planning. In lack of public cost-related datasets, we further present OpenCATP, the first platform for cost-aware planning evaluation. Experiments on OpenCATP show that CATP-LLM outperforms GPT-4 even when using Llama2-7B as its backbone, with the average improvement of 28.2%-30.2% higher plan performance and 24.7%-45.8% lower costs even on the challenging planning tasks. The codes and dataset will be available at: https://github.com/duowuyms/OpenCATP-LLM
Figure 1: We introduce CATP-LLM, an efficient framework that enables LLMs to create complex tool plans for challenging tasks while striking a good balance between plan performance and execution costs. We further present OpenCATP, a cost-aware evaluation platform with the Quality of Plan (QoP) metric to quantitatively evaluate a plan based on its task performance and execution costs (e.g. time and memory consumption). We demonstrate that CATP-LLM with the small Llama2-7B as backbone significantly outperforms GPT-powered methods in terms of higher QoP, with significantly lower execution costs and higher or comparable task performance.
@@ 1 Introduction
The recent advancement in large language models (LLMs) [1, 11] have highlighted their extraordinary ability in tool planning [35, 47, 9], where LLMs can schedule and execute external tools (e.g. vision models [22, 37] and python programs [12, 15]) to tackle complex tasks based on task descriptions. For instance, when confronted with the task “How to generate the caption of a low-resolution image in German”, the LLM may devise the following tool plan: (1) use a super resolution tool to enlarge the image; (2) employ the image captioning model to generate caption; (3) execute a translator to convert the caption into German. Due to the flexibility of tool usage and composition, this paradigm has shown exceptional performance, especially in solving complex tasks including visual question answering [40, 48], image editing [49, 44, 25] and mathematical reasoning [51, 46], paving the way for the development of more general AI systems [12, 10]. Furthermore, it offers significant practical prospects by effectively utilizing existing tools and models, thereby reducing the need for developing and training new models [37].
Despite promising, using LLMs for tool planning still faces practical issues related to the plan execution costs. Specifically, the execution of tools will inevitably cause computation costs such as execution time and memory consumption. This makes it essential for the LLMs to carefully take the execution overhead of each tool into account when devising tool plans for specific tasks. Otherwise, they may generate expensive tool plans of which the execution costs outweigh task performance, ultimately hindering their practical applications (e.g. Figure 1③).
@@
  • First, how to enable LLMs for non-sequential tool planning? Non-sequential planning refers to the capability to create nonlinear plans that comprise multiple branches for concurrent tool execution. It not only facilitates more efficient parallel processing of tools to reduce execution costs (e.g. execution time), but also enhances the ability of LLMs to solve complicated tasks that require complex tool plans [10], as exemplified in Figure 1④. Nevertheless, implementing non-sequential planning is challenging. Recent insights have revealed that it can be difficult for LLMs to understand and generate nonlinear-structured contents through natural language prompts [16, 32, 50, 4], not to mention the complex tool dependencies that must be strictly followed during the generation of structured plans.
@@
  • Second, how to enable LLMs to achieve the optimal performance-cost trade-off in tool planning? The key obstacle lies in the conflicts between plan performance and execution costs. Generally, simple plans often incur lower costs but may compromise performance, while complex plans can enhance performance but at the expense of higher costs. Moreover, the costs of executing tool plans can be affected by the input data associated with the task, which further complicates the problem. Consequently, striking a good balance between performance and costs in tool planning remains a significant challenge for LLMs, which requires a deep understanding of the performance-cost trade-off and the impacts of input data on execution costs.
@@
  • Last but not least, in lack of public datasets that consider tool execution costs, how to evaluate the effectiveness of LLMs in cost-aware tool planning?
@@
Prior studies [22, 28, 37, 10] fail to achieve cost-aware tool planning as most of them only support sequential planning and overlook the tool execution costs. To fill this research gap, in this paper, we propose the Cost-Aware Tool Planning with LLMs (CATP-LLM) framework, which for the first time provides a coherent design to efficiently enable LLMs for cost-aware tool planning. Specifically, we propose the following techniques to systematically overcome the aforementioned challenges:
@@
  • First, we design a tool planning language (TPL) to enhance the non-sequential planning capability of LLMs. TPL transforms tools and their dependencies into learnable tokens, which, like language tokens, can be trained to learn the optimal representations for LLMs to understand their semantic information. TPL then encodes any tool plan, including non-sequential ones, as a sequence of tool and dependency tokens, so that the LLMs can generate diverse tool plans by simply predicting these tokens. Besides, it also introduces some special tokens to describe the structural information of tool plans to facilitate LLMs to generate tool plans of complex structures.
@@
  • Based on the TPL, we then propose an efficient cost-aware offline reinforcement learning (CAORL) algorithm to address the second challenge. Specifically, CAORL introduces a context augmentation scheme for LLMs to capture the impacts of input data on the tool execution costs for effective cost-aware planning. It then fine-tunes LLMs based on the efficient offline RL algorithm [2, 7, 6] with a dedicated reward model to guide the LLM to minimize plan execution costs while maximizing plan performance. Besides, a data generation method is also designed to generate the training data for fine-tuning.
@@
  • To comprehensively assess the efficacy of LLMs in cost-aware tool planning, we establish a new platform OpenCATP. The key features of OpenCATP lie in three aspects: (1) it allows the measurement of plan costs; (2) it incorporates a unified metric Quality of Plan (QoP) to quantitatively evaluate the plan quality based on its performance and execution costs; (3) it introduces more complex tasks for the non-sequential planning evaluation (e.g. Figure 1).
@@
We compare CATP-LLM with existing methods on OpenCATP. Results demonstrate that CATP-LLM outperforms GPT-3.5 [29] and GPT-4 [1] with Llama2-7B [41] as the backbone LLM. To be specific, it achieves 1.15x-1.6x and 2.29x-9.13x higher QoP in the case of sequential planning and non-sequential planning, respectively, with lower execution costs and higher or comparable plan performance.
@@
  • We develop a new platform OpenCATP for the research community to comprehensively assess the performance of LLMs in cost-aware tool planning.
@@ 2.1 Tool Planning with LLMs
Large language models (LLMs) [11, 41] have shown exceptional performance in understanding task descriptions and deriving tool execution plans based on the provided toolkit to address the target tasks [8, 26, 40, 24]. Therefore, tool planning with LLMs has emerged as a promising avenue toward developing general AI systems [12, 10]. In pursuit of this vision, numerous works have been proposed to enhance the tool planning abilities of LLMs, which can be categorized into two paradigms. The first paradigm leverages prompt engineering [5, 34, 23] to guide LLMs for tool scheduling, which has been recently extended by adding in-context demonstrations [49, 30, 40, 12], introducing more tools [28, 44], or designing more dedicated planning procedures [37, 26, 22]. Another line of work involves fine-tuning LLMs for tool planning through instruction-tuning [47, 35, 13, 31]. Recent advancements have also introduced reinforcement learning techniques [10, 33] to further augment LLMs for tool planning.
In Table 1, we compare CATP-LLM with other LLM-based tool planning methods. Existing methods exhibit limited support for non-sequential planning, restricting the full potential of LLMs in cost-aware tool planning. For instance, prior work [22] combines formal language with natural language to improve the planning ability of the LLM, but it necessitates domain experts to handcraft specialized prompts, which can be labor-intensive. Although methods proposed in [37, 26] integrate non-sequential plans in JSON-like format into prompts as demonstrations, recent insights [16, 32, 39, 50] suggest that LLMs may not effectively understand and generate nonlinear-structured information through language instructions, thereby hindering these methods to function effectively. In contrast, CATP-LLM incorporates a tool planning language to empower LLMs for non-sequential planning to effectively generate tool plans of various structures. Moreover, while existing methods overlook the execution costs of tool plans, CATP-LLM introduces an efficient algorithm to fine-tune LLMs for cost optimization.
@@ 2.2 Reinforcement Learning for Tool Planning
Reinforcement learning (RL) [43, 33] has been proposed to enhance LLMs for tool planning, which enables LLMs to understand their actions and adjust behavior accordingly by replaying plan execution feedback to LLMs. However, existing methods [10, 33] necessitate completing plan execution before replaying feedback to LLMs, resulting in coarse-grained planning. In contrast, CATP-LLM introduces an efficient RL algorithm tailored for cost-aware tool planning. It allows LLMs to access the plan quality each time a decision is made (e.g. adding a new tool to the plan). Such fine-grained perception empowers LLMs for more effective cost-aware planning (e.g. selecting a cheaper tool if the current plan costs are already excessive).
@@ 2.3 Datasets for Tool Planning
With the advancement of tool planning with LLMs, numerous datasets have been established to access LLMs’ capability in scheduling tools for task solving [42, 38, 27, 54, 21]. From this line of work, OpenAGI [10] is the most relevant to this paper, which offers a quantitative scheme to objectively evaluate the performance of LLM-generated tool plans in various tasks. However, existing works, including OpenAGI, overlook the tool execution costs, the crucial aspect that significantly affects the efficacy of tool planning in real-world scenarios. To bridge this gap, this work develops a new platform OpenCATP that for the first time integrates tool cost information and introduces a comprehensive scheme for the evaluation of cost-aware tool planning.
@@ 3 CATP-LLM Design
  • Tool planning language (TPL). TPL encodes tools and their dependencies into trainable tokens, and represents any tool plan as a token sequence augmented with special tokens to capture the plan structural information (§$\S$3.1).
@@ 3.1 Tool Planning Language
Given a set of tools $\mathcal{T}=\{t_{1},t_{2},\cdots\}$ and their dependencies $\mathcal{D}=\{d_{1},d_{2},\cdots\}$, a tool plan can be expressed as a directed acyclic graph (DAG) $p=(\{t_{i}\}_{i=1}^{n},\{d_{j}\}_{j=1}^{m})$, where node $t_{i}\in\mathcal{T}$ indicates a specific tool and edge $d_{j}\in\mathcal{D}$ denotes a specific dependency (e.g. a resource dependency where one tool accepts the outputs of another tool as inputs). To enable the LLM to generate tool plans of various topological structures, the proposed TPL comprises the following key designs.
@@
Anything as token. Inspired by the work in [13], we formulate tools and dependencies as tokens [^1] The work in [13] also represents tools as tokens. Our proposed TPL distinguishes itself from [13] by extending this idea to tool dependencies and explicitly formulating a tool plan as a sequence of tool and dependency tokens (explained later in this section)., denoted as $[\mathcal{T}]=\{[t_{1}],[t_{2}],\cdots\}$ and $\langle\mathcal{D}\rangle=\{\langle d_{1}\rangle,\langle d_{2}\rangle,\cdots\}$, respectively. Similar to regular word tokens, each tool or dependency token is parameterized by an embedding vector that can be trained to learn the optimal representations for the LLM to understand its semantic information. The key benefit of this approach is that it can accommodate new tools or dependencies by simply adding new tokens, allowing the LLM to support massive tools and dependencies.
@@
where $[t_{i}]\in[\mathcal{T}]$ denotes the $i$-th tool token, and $\langle d_{j}^{i}\rangle\in\langle\mathcal{D}\rangle$ represents the dependency between the $i$-th tool and the previous tool. In our practical implementation, we introduce an additional token $\langle task\rangle$ to indicate a special dependency that a tool accepts the data provided by the task (e.g. an image) as inputs. In particular, $[SoP]$, $[EoP]$ and $\langle SoD\rangle$, $\langle EoP\rangle$ are special structure tokens indicating the start, end of the plan and start, end of the dependencies associated with a tool, respectively. These structure tokens are used to facilitate the LLM to effectively perceive the plan structural information. Through such a flexible representation, our TPL can represent tool plans of diverse structures, enhancing the non-sequential planning ability of the LLM. Figure 3 exemplifies the use of TPL for encoding various tool plans.

### v3

@@ Abstract
Utilizing large language models (LLMs) for tool planning has emerged as a promising avenue for developing general AI systems, where LLMs automatically schedule external tools (e.g., vision models) to tackle complex tasks based on task descriptions. To push this paradigm toward practical applications, it is crucial for LLMs to consider tool execution costs (e.g., execution time) for tool planning. Unfortunately, prior studies overlook the tool execution costs, leading to the generation of expensive plans whose costs outweigh their benefits in terms of task performance. To fill this gap, we propose the Cost-Aware Tool Planning with LLMs (CATP-LLM) framework, which for the first time provides a coherent design to empower LLMs for cost-aware tool planning. Specifically, To facilitate efficient concurrent tool execution and cost reduction, we design a tool planning language to enhance the LLM for creating multi-branch non-sequential plans. Moreover, we propose a cost-aware offline reinforcement learning algorithm to fine-tune the LLM to optimize the performance-cost trade-off in tool planning. In the lack of public cost-related datasets, we further present OpenCATP, the first dataset for cost-aware planning, which comprises 11,100 evaluation samples from diverse tasks. Extensive experiments show that CATP-LLM outperforms GPT-4 even when using Llama2-7B as its backbone, with the average improvement of 1.5%-93.9% in terms of plan quality. Codes and dataset are available at: https://github.com/duowuyms/OpenCATP-LLM.
Figure 1: We introduce CATP-LLM, an efficient framework that enables LLMs to create logically complex tool plans for task solving while striking a good balance between plan performance and tool execution costs. We further present OpenCATP, a cost-aware evaluation dataset with the Quality of Plan (QoP) metric to quantitatively evaluate a plan based on its task performance and execution costs (e.g., time and memory consumption). We demonstrate that CATP-LLM with the small Llama2-7B as backbone significantly outperforms GPT-powered methods in terms of higher QoP, with significantly lower execution costs and higher or comparable task performance.
@@ 1 Introduction
Recent advances in large language models (LLMs) [1, 16] have highlighted their extraordinary ability in tool planning [43, 55, 14], where LLMs can schedule and execute external tools (e.g., vision models [28, 45] and Python programs [17, 20]) to tackle complex tasks based on task descriptions. For instance, when confronted with the task “How to generate the caption of a low-resolution image in German”, the LLM may devise the following tool plan: (1) use a super resolution tool to enlarge the image; (2) employ the image captioning model to generate caption; (3) execute a translator to convert the caption into German. Due to the flexibility of tool usage and composition, this paradigm has shown exceptional performance, especially in solving complex tasks including visual question answering [48, 56], image editing [57, 52, 31] and mathematical reasoning [60, 54], paving the way for the development of more general AI systems [17, 15]. Furthermore, it offers significant practical prospects by effectively utilizing existing tools and models, thereby reducing the requirement for developing and training new models [45].
Despite its promising potential, using LLMs for tool planning still faces practical issues related to plan execution costs. Specifically, the execution of tools will inevitably cause computation costs such as execution time and memory consumption, as depicted in Figure 1. This makes it essential for LLMs to carefully take into account tool execution overhead when devising tool plans for specific tasks. Otherwise, they may generate expensive tool plans whose execution costs outweigh task performance, ultimately hindering their practical applications (e.g., Figure 1③).
@@
  • First, how to enable LLMs for non-sequential tool planning for potential cost reduction? Non-sequential planning refers to the capability to create logically nonlinear plans that comprise multiple branches for concurrent tool execution. It not only facilitates more efficient parallel processing of tools to reduce execution costs (e.g., execution time), but also enhances the ability of LLMs to solve complicated tasks that require complex tool plans [15], as exemplified in Figure 1④. Nevertheless, implementing non-sequential planning remains a challenge as recent insights have revealed that it can be difficult for LLMs to understand and generate logically nonlinear content through natural language prompts [21, 39, 58, 5].
@@
  • Second, how to enable LLMs to achieve the optimal performance-cost trade-off in tool planning? The key obstacle lies in the conflicts between plan performance and execution costs. Generally, simple plans often incur lower costs but may compromise performance, while complex plans can enhance performance but at the expense of higher costs. In addition, the costs of executing tool plans can be affected by the input data associated with the task (e.g., larger images increase object detection overhead). Consequently, striking a good balance between performance and costs in tool planning remains a significant challenge for LLMs, which requires a deep understanding of the performance-cost trade-off and the impacts of input data on execution costs.
@@
  • Finally, in the lack of public datasets that consider non-sequential planning and tool execution costs, how to evaluate the efficacy of LLMs in cost-aware tool planning?
@@
Prior studies [28, 35, 45, 15] fail to achieve cost-aware tool planning, as most of them only support sequential planning and overlook the tool execution costs. To fill this research gap, in this paper, we propose the Cost-Aware Tool Planning with LLMs (CATP-LLM) framework, which for the first time provides a coherent design to efficiently enable LLMs for cost-aware tool planning. Specifically, we propose the following techniques to systematically overcome the aforementioned challenges:
@@
  • First, we design a tool planning language (TPL) to enhance the non-sequential planning capability of LLMs. TPL transforms tools and their dependencies into learnable tokens, which, like language tokens, can be trained to learn the optimal representations for LLMs to understand their semantic information. TPL then encodes any tool plan, including non-sequential ones, as a sequence of tool and dependency tokens, so that the LLMs can generate diverse tool plans simply by predicting these tokens. Additionally, it also introduces some special tokens to describe the structural information of tool plans to facilitate LLMs to generate tool plans of complex structures.
@@
  • Based on the TPL, we then propose an efficient cost-aware offline reinforcement learning (CAORL) algorithm to address the second challenge. Specifically, CAORL introduces a context augmentation scheme for LLMs to capture the impacts of input data on the tool execution costs for effective cost-aware planning. It then fine-tunes LLMs based on the efficient offline RL algorithm [2, 12, 8] with a dedicated reward model to guide the LLM to minimize plan execution costs while maximizing plan performance. Besides, a data generation method is also designed to generate the training data for the LLM fine-tuning.
@@
  • To comprehensively assess the efficacy of LLMs in cost-aware tool planning, we establish a new dataset OpenCATP. The key features of OpenCATP lie in three aspects: (1) it allows the measurement of plan costs; (2) it incorporates a unified metric Quality of Plan (QoP) to quantitatively evaluate the plan quality based on its performance and execution costs; (3) it comprises 11,100 samples including 2,400 samples for the non-sequential planning evaluation (e.g., Figure 1).
@@
We compare CATP-LLM with existing methods on OpenCATP. Results demonstrate that CATP-LLM outperforms GPT-3.5 [36] and GPT-4 [1] with Llama2-7B [49] as the backbone LLM. To be specific, it achieves 1.02x-1.41x and 1.92x-5.99x higher QoP in the case of sequential planning and non-sequential planning, respectively, with lower execution costs and higher or comparable plan performance.
@@
  • We develop a new dataset OpenCATP for the research community to comprehensively assess the performance of LLMs in cost-aware tool planning.
@@ 2.1 Tool Planning with LLMs
Large language models (LLMs) [16, 49] have shown exceptional performance in understanding task descriptions and deriving tool execution plans based on the provided toolkit to address the target tasks [13, 32, 48, 30]. Therefore, tool planning with LLMs has emerged as a promising avenue to develop general AI systems [17, 15]. In pursuit of this vision, numerous works have been proposed to enhance the tool planning abilities of LLMs, which can be categorized into two paradigms. The first paradigm leverages prompt engineering [7, 42, 29] to guide LLMs for tool scheduling, which has recently been extended by adding in-context demonstrations [57, 37, 48, 17], introducing more tools [35, 52], or designing more dedicated planning procedures [45, 32, 28, 24]. Another line of work involves fine-tuning LLMs for tool planning through instruction tuning [55, 43, 18, 38]. Recent advancements have also introduced reinforcement learning techniques [15, 40] to further augment LLMs for tool planning.
In Table 1, we compare CATP-LLM with other LLM-based tool planning methods. Existing methods exhibit limited support for non-sequential planning, restricting the full potential of LLMs in cost-aware tool planning. For instance, prior work [28] combines formal language with natural language to improve the planning ability of the LLM, but it necessitates domain experts to handcraft specialized prompts, which can be labor-intensive. Although methods proposed in [45, 32] integrate non-sequential plans in JSON-like format into prompts as demonstrations, recent insights [21, 39, 47, 58] suggest that LLMs may not effectively understand and generate logically nonlinear information through language instructions, thereby hindering these methods to function effectively. In contrast, CATP-LLM incorporates a tool planning language to empower LLMs for non-sequential planning to effectively generate tool plans of various structures. Moreover, while existing methods overlook tool execution costs, CATP-LLM introduces an efficient algorithm to fine-tune LLMs for cost optimization.
@@ 2.2 Reinforcement Learning for Tool Planning
Reinforcement learning (RL) [51, 40] has been proposed to enhance LLMs for tool planning, which enables LLMs to understand their actions and adjust behavior accordingly by replaying plan execution feedback to LLMs. However, existing methods [15, 40] suffer from coarse-grained planning, as they replay feedback to LLMs only at the completion of plan execution without any intermediate guiding signals. In contrast, CATP-LLM introduces an efficient RL algorithm tailored for cost-aware tool planning. It allows LLMs to access the plan quality each time a decision is made (e.g., adding a new tool to the plan). Such fine-grained perception empowers LLMs for more effective cost-aware planning (e.g., selecting a cheaper tool if the current plan costs are already excessive).
@@ 2.3 Datasets for Tool Planning
With the advancement of tool planning with LLMs, numerous datasets have been established to access LLMs’ capability in scheduling tools for task solving [50, 46, 33, 64, 15]. However, existing datasets overlook non-sequential planning and tool execution costs, the crucial aspects that significantly affect the efficacy of tool planning in real-world scenarios. To bridge this gap, this work develops a new dataset OpenCATP that for the first time introduces many challenging tasks for non-sequential planning evaluation and a comprehensive scheme to evaluate plan quality based on its task performance and execution costs.
@@ 3 CATP-LLM Design
  • Tool planning language (TPL). TPL encodes tools and their dependencies into trainable tokens, and represents any tool plan as a sequence of tokens augmented with special tokens to capture the structural information of the plan (§$\S$3.1).
@@ 3.1 Tool Planning Language
Given a set of tools $\mathcal{T}=\{t_{1},t_{2},\cdots\}$ and their dependencies $\mathcal{D}=\{d_{1},d_{2},\cdots\}$, a tool plan can be expressed as a directed acyclic graph (DAG) $p=(\{t_{i}\}_{i=1}^{n},\{d_{j}\}_{j=1}^{m})$, where node $t_{i}\in\mathcal{T}$ indicates a specific tool and edge $d_{j}\in\mathcal{D}$ denotes a specific dependency (e.g., a resource dependency where one tool accepts the outputs of another tool as inputs). To enable the LLM to generate tool plans of various topological structures, the proposed TPL comprises the following key designs.
@@
Anything as token. Inspired by the work in [18], we formulate tools and dependencies as tokens [^1] The work in [18] also represents tools as tokens. Our proposed TPL distinguishes itself largely from [18] by representing tool dependencies as tokens and explicitly formulating a tool plan as a sequence of tool and dependency tokens (explained later in this section)., denoted as $[\mathcal{T}]=\{[t_{1}],[t_{2}],\cdots\}$ and $\langle\mathcal{D}\rangle=\{\langle d_{1}\rangle,\langle d_{2}\rangle,\cdots\}$, respectively. Similar to regular word tokens, each tool or dependency token is parameterized by an embedding vector that can be trained to learn the optimal representations for the LLM to understand its semantic information. The key benefit of this approach is that it can accommodate new tools or dependencies simply by adding new tokens, allowing the LLM to support massive tools and dependencies.
@@
where $[t_{i}]\in[\mathcal{T}]$ denotes the $i$-th tool token, and $\langle d_{j}^{i}\rangle\in\langle\mathcal{D}\rangle$ represents the dependency between the $i$-th tool and the previous tool. In our practical implementation, we introduce an additional token $\langle task\rangle$ to indicate a special dependency that a tool accepts the data provided by the task (e.g., an image) as inputs. In particular, $[SoP]$, $[EoP]$ and $\langle SoD\rangle$, $\langle EoP\rangle$ are special structure tokens indicating the start, end of the plan and start, end of the dependencies associated with a tool, respectively. These structure tokens are used to facilitate the LLM to effectively perceive the plan structural information. Through such a flexible representation, our TPL can represent tool plans of diverse structures, enhancing the non-sequential planning ability of the LLM. Figure 3 exemplifies the use of TPL for encoding various tool plans.